# Large Intelligent Surfaces with Low-End Receivers: From Scaling to Antenna and Panel Selection

### v2

@@ VI Numerical Results
In this section, we provide numerical examples to gain more insights about the aforementioned methods and derivations. As performance metric, we consider a lower bound on the spectral efficiency given by $R_{\text{a}}=\text{log}_{2}(1+\gamma)$, which corresponds to an achievable rate. This achievable rate comes from assuming a Gaussian additive distortion term in the Bussgang decomposition, corresponding to the worst case scenario [31]. The UE is assumed to be at distance $d=25\lambda$ from the center of LIS transmitting with transmit power $P$ such that SNR$=10dB$ at the center of LIS. The gray areas in all the figures in this section illustrate the range of parameteres for which the UE is in far-field of the LIS. For the hardware distortion model, we use the measurements from one of the 3GPP reports [25], which also estimates the parameters for the memory-less polynomial model (4) based on real-world RF hardware at different frequencies and bandwidths. To have a better understanding of the difference in hardware distortion effects for different RF components, we have employed (13) and (14) to calculate $\tilde{g}$ and $\kappa$ based on a range of measurements in [25], which have been performed at the frequencies $2.1$ GHz and $28$ GHz, for GaA, CMOS, and GaN amplifiers. Table I summarizes the results. Note that, for ideal hardware, we would have $|\tilde{g}|^{2}=1$ and $\kappa=0$.
@@ VI-A Scaling Analysis Example
We begin with studying the scaling performance of LISs for different levels of hardware distortion. Let us consider the model based on a Gallium Nitride (GaN) amplifier designed for operation at 2.1 GHz as a case study, where the model parameters have been estimated from input-output measurements at a sample rate of 200 MHz and a signal bandwidth of 40 MHz. We would have similar results and conclusions if we select another data set from [25] and table I. The report provides the coefficients for normalized inputs, i.e., $\hat{a}_{2k+1}$ in (11) and (12), for 3rd, 5th, 7th, and 9th order non-linearity. We consider a discrete LIS with antenna elements separated by $\lambda/2$ on a rectangular grid.
@@
In Fig. 4 we consider the case with perfect per-antenna AGC, where we plot the achievable rate for different levels of back-off. So as to evaluate the approximation error for the proposed close-form expressions, we consider both the closed-form expression and the exact numerical values. Firstly, we can see that the close-form approximations overlap closely with the numerical values, which means the approximations are valid and can be used for analytical results with very low approximation error. Secondly, we can see that similar to Fig. 3 which was without per-antenna perfect AGC, the hardware distortion degrades the system performance significantly, which implies that hardware distortion may potentially limit the performance in LIS scenarios no matter if we have perfect AGC or not.
@@ VI-B Antenna and Panel selection
In Fig. 5, the performance of using the proposed optimal antenna selection in comparison to using the dominant antennas, i.e. antennas with highest received power, is illustrated. We have used the 9th order distortion model for a GaN amplifier at $2.1$ GHz, as described in the previous section, with a $7$ dB of back-off. The UE transmit power is again selected such that the SNR at the center of LIS reaches $10$ dB, which corresponds to the respective value of $P_{\text{max}}$ in (11). We also assume that $N_{\text{max}}=\lceil 0.1N\rceil$ is selected as the antenna selection constraint. As we can see in Fig. 5, adopting the optimal antenna selection can improve the system performance significantly for medium to large LIS radius. We can also see that the gain from adopting antenna selection is negligible if the UE is in the far-field of the LIS.
@@
In Fig. 8, we have evaluated the performance of the approximate closed-form solution to the panel selection problem in comparison to the numerical optimal solution and the dominant panel selection. We have used the 3rd order non-linearity model for the hardware distortion as motivated in section V. The panel selection constraint and panel placement is the same as in Fig. 6. As we see, the optimal panel selection and closed-form panel selection outperform the dominant panel selection significantly, and the close-form panel selection performance is very close to the numerical optimal method, which implies that the proposed close-form solution is accurate enough to use in LIS panel selection problems with hardware distortion. In Fig. 9, we compare the performance gain of adopting such schemes for 3 different distance of the UE. We can see that the gain is higher if the UE is more into the near-field of the LIS, and it is still significant if the UE is placed at further distances. The results from Fig. 5, 6, 8, and 9 show that we can improve the system performance by performing the proposed antenna and panel selection without using more resources in terms of extra antennas and panels. This implies that the proposed selection methods can improve the system energy efficiency (EE) which is favorable when scaling up LIS systems for future generations of wireless networks.
@@ Appendix A Proof of Lemmas
[…]
@@ Appendix B Algorithms
Algorithm 1 provides a simple low-complexity numerical method to solve (31), where the accuracy of the solution depends on the number of iterations, $T$.
@@
[…]

### v3

@@ VI Numerical Results
In this section, we provide numerical examples to gain more insights about the aforementioned methods and derivations. As performance metric, we consider a lower bound on the spectral efficiency given by $R_{\text{a}}=\text{log}_{2}(1+\gamma)$, which corresponds to an achievable rate. This achievable rate comes from assuming a Gaussian additive distortion term in the Bussgang decomposition, corresponding to the worst case scenario [37]. The UE is assumed to be at distance $d=25\lambda$ from the center of LIS transmitting with transmit power $P$ such that SNR=10dB=10dB at the center of LIS. The gray areas in all the figures in this section illustrate the range of parameters for which the UE is in far-field of the LIS. For the hardware distortion model, we use the measurements from one of the 3GPP reports [29], which also estimates the parameters for the memory-less polynomial model (4) based on real-world RF hardware at different frequencies and bandwidths. To have a better understanding of the difference in hardware distortion effects for different RF components, we have employed (13) and (14) to calculate $\tilde{g}$ and $\kappa$ based on a range of measurements in [29], which have been performed at the frequencies $2.1$ GHz and $28$ GHz, for GaA, CMOS, and GaN amplifiers. Table I summarizes the results. Note that, for ideal hardware, we would have $|\tilde{g}|^{2}=1$ and $\kappa=0$.
@@ VI-A Scaling Analysis Example
We begin with studying the scaling performance of LISs for different levels of hardware distortion. Let us consider the model based on a Gallium Nitride (GaN) amplifier designed for operation at 2.1 GHz as a case study, where the model parameters have been estimated from input-output measurements at a sample rate of 200 MHz and a signal bandwidth of 40 MHz. We would have similar results and conclusions if we select another data set from [29] and table I. The report provides the coefficients for normalized inputs, i.e., $\hat{a}_{2k+1}$ in (11) and (12), for 3rd, 5th, 7th, and 9th order non-linearity. We consider a discrete LIS with antenna elements separated by $\lambda/2$ on a rectangular grid. Selecting this value for antenna-spacing is based on the fact that $\lambda/2$ spacing allows capturing the full spatial degrees of freedom, while facilitating practical implementation [1, 6]. Nevertheless, it can be verified that under the considered framework, the same conclusions can be derived for antenna spacings below $\lambda/2$.
@@
In Fig. 4 we consider the case with perfect per-antenna AGC, where we plot the achievable rate for different levels of back-off. So as to evaluate the approximation error for the proposed close-form expressions, we consider both the closed-form expression and the exact numerical values. Firstly, we can see that the close-form approximations from lemma III.1 overlap closely with the numerical exact values, which means the approximations are valid and can be used for analytical results with very low approximation error. As a result, we can approximate problem (42) with (45) with a negligible loss in performance. Secondly, we can see that similar to Fig. 3, which was without per-antenna perfect AGC, the hardware distortion degrades the system performance significantly, which implies that hardware distortion may potentially limit the performance in LIS scenarios no matter if we have perfect AGC or not.
@@ VI-B Antenna and Panel selection
In Fig. 5, the performance of using the proposed optimal antenna selection in comparison to using the dominant antennas, i.e., antennas with highest received power, is illustrated. Note that, in the single-UE scenario without hardware distortion, dominant antenna selection is the optimal selection approach in terms of capacity, while for multi-UE scenarios, the loss with respect to optimal sum-rate maximization is marginal for large enough number of RX-chains [38]. We have used the 9th order distortion model for a GaN amplifier at $2.1$ GHz, as described in the previous section, with a $7$ dB of back-off. The UE transmit power is again selected such that the SNR at the center of LIS reaches $10$ dB, which corresponds to the respective value of $P_{\text{max}}$ in (11). We also assume that $N_{\text{max}}=\lceil 0.1N\rceil$ is selected as the antenna selection constraint. As we can see in Fig. 5, adopting the optimal antenna selection can improve the system performance significantly for medium to large LIS radius. We can also see that the gain from adopting antenna selection is negligible if the UE is in the far-field of the LIS.
@@
In Fig. 8, we have evaluated the performance of the approximate closed-form solution to the panel selection problem in comparison to the numerical optimal solution and the dominant panel selection. We have used the 3rd order non-linearity model for the hardware distortion as motivated in section V. The panel selection constraint and panel placement is the same as in Fig. 6. As we see, the optimal panel selection and closed-form panel selection outperform the dominant panel selection significantly, and the close-form panel selection performance is very close to the numerical optimal method, which implies that the proposed close-form solution is accurate enough to use in LIS panel selection problems with hardware distortion. In Fig. 9, we compare the performance gain of adopting such schemes for 3 different distance of the UE. We can see that the gain is higher if the UE is more into the near-field of the LIS, and it is still significant if the UE is placed at further distances. The results from Fig. 5, 6, 8, and 9 show that we can improve the system performance by performing the proposed antenna and panel selection without using more resources in terms of extra antennas and panels. This implies that the proposed selection methods can improve the system EE which is favorable when scaling up LIS systems for future generations of wireless networks.
@@
VI-C Distortion Correlation
As remarked in Section III, the distortion covariance matrix $\bm{C_{\eta\eta}}$ is in general non-diagonal, and assuming it diagonal, i.e., disregarding the correlation, results in an over-estimation of the SNDR. In this section, we study the impact of disregarding the distortion correlation on the scaling behavior and on the performance of the panel selection schemes. Assuming a 3rd order non-linearity model, i.e. $L=2$, for the non-linearity model in (4), the non-diagonal distortion covariance matrix elements are given by
$[\bm{C_{\eta\eta}}]_{ij}=2a_{3,i}\bar{a}_{3,j}|\rho_{ij}|^{2}\rho_{ij},$ (56)
where $\rho_{ij}=\mathbb{E}\{x_{i}\bar{x}_{j}\}$, and $x_{i}$ is the input signal to the $i$’th antenna. While characterizing the non-diagonal elements of $\bm{C_{\eta\eta}}$ for higher order non-linearities is beyond the scope of this paper, we expect the 3rd order case to capture the main impact of disregarding distortion correlation.
In Fig. 11, we have analyzed the scaling behavior of LIS by comparing the achievable rate $R_{\text{a}}$ for different levels of back-off, both with and without considering the distortion correlation. The transmit power is selected such that $\text{SNR}=0\text{ dB}$ at the center of the LIS. We can see that as we scale up the LIS, the impact of disregarding the distortion grows progressively. However, for reasonably large LISs, the impact is still quite limited. On the other hand, taking into account the distortion correlation will only strengthen the important finding that hardware distortion has a significant impact on the performance of scaled-up LIS systems.
In Fig. 12 we analyze the performance of the panel selection schemes when these are either aware or unaware of the distortion correlation. For that purpose, we have plotted the achievable data rate with hardware distortion correlation for three cases: dominant panel selection, sub-optimal panel selection which performs the panel selection unaware of the distortion correlation, and the optimal panel selection which takes into account the distortion correlation for panel selection. All the simulation parameters are the same as in Fig. 8. We can see that, although Fig. 11 shows that disregarding distortion correlation can result in considerable over-estimation of the data rate, this approximation can be leveraged for low-complexity panel selection, still resulting in close-to-optimum performance and significant gain compared to the baseline approach.
Figure 11: Achievable data rate vs LIS radius in terms of $\lambda$, with and without considering distortion correlation. The UE is at distance $d=25\lambda$ from the center of LIS.
Figure 12: Achievable data rate with distortion correlation vs number of panels. The UE is at distance $d=25\lambda$ from the center of LIS, the panel selection constraint in (47) is $N_{\text{max}}=\lceil 0.1N\rceil$, and Each panel consists of $M=16$ antennas with $\lambda/2$ spacing. The distance between the centers of adjacent panels is $\delta_{p}=5\lambda$.
@@ Appendix A Proof of Lemmas
$\int\frac{1}{(d^{2}+x)^{\frac{3}{2}}}dx=\frac{-2}{\sqrt{d^{2}+x}},$ (64)
$\int\frac{1}{(d^{2}+x)^{3}}dx=\frac{-\frac{1}{2}}{(d^{2}+x)^{2}},$ (65)
we can find the results in (24) and (25), and a simple factorization gives (26).
@@ Appendix B Algorithms
In this Section, we provide algorithms to showcase that the optimization problems formulated in the paper can be solved numerically with limited computation complexity. Algorithm 1 provides a simple low-complexity numerical method to solve (32), where the accuracy of the solution depends on the number of iterations, $T$.
@@
 Define:  $R_{\mathrm{ub}}=$ (38),  $R_{\mathrm{lb}}=$ (39),  $\Delta=R_{\mathrm{ub}}-R_{\mathrm{lb}}$$R_{\text{opt}}=R_{\mathrm{lb}}$$R_{\text{new}}=R_{\mathrm{lb}}$$T\in\mathcal{N}$ where $T>>1$.